# Life Finds a Way: Emergence of Cooperative Structures in Adaptive Threshold Networks

### v3

@@ Abstract
We present a random threshold-directed network model [24] that integrates node-specific traits with dynamic edge formation and node removal, simulating arbitrary levels of cooperation and competition. In our framework, intrinsic node values determine directed links through various threshold rules. Our model generates a multi-digraph with signed edges (reflecting support/antagonism, labeled “help”/“harm”), which ultimately yields two parallel yet interdependent threshold graphs. Incorporating temporal growth and node turnover in our approach allows exploration of the evolution, adaptation, and potential collapse of communities and reveals phase transitions in both connectivity and resilience.
@@ 2 Model
2.1 Description of Vanilla Model
Using the python library Networkx, we dynamically generated a Directed Multi-Graph $G=(V,E)$ where $V$ is the set of vertices (or nodes) and $E$ is the set of edges (or links). At each discrete time-step the model adds one new vertex to the digraph and endows it with four independent traits that represent a subset of its causal properties to give or receive help and harm. The newcomer then “tests” every pre-existing vertex in both directions. For each ordered pair, the absolute difference between the relevant traits is compared with a random Poisson-scaled threshold: if the gap is small enough, then an edge is created. This threshold is also modulated by a global help-to-harm ratio, and a small random multiplicity is allowed, so a single pair can accumulate several parallel edges (we set a maximum of 6 to calculations, but this can be altered) to represent multiple affordances or adverse interactions.
@@ 2.1 Description of Vanilla Model
Figure 1: Sample Graph Generation
@@
Step 2: Node B added, harm (red) edge added $B\to A$ as $|A_{\gamma_{in}}-B_{\gamma_{out}}|<P_{AB}$, and help (blue) edge added $A\to B$ as $|A_{h_{out}}-B_{h_{in}}|<P_{AB}$.
@@
Once edges are formed, every vertex ages by one tick. A node whose age now exceeds its prescribed lifespan is removed—and all of its incident edges with it—under either of two conditions: it has no incoming edges at all, or more than a fixed fraction (usually 50 %) of its incoming edges are harmful. If there are enough helpful edges, then nodes survive indefinitely. Edges never disappear on their own; they vanish only when one of their endpoints is deleted. Repeating this birth–binding–culling cycle yields the evolving signed network on which the subsequent experiments are run.
AABBCCDDEE
Figure 2: Death Criterion
Death threshold ($\eta$) passed. Node C (and all its incoming and outgoing edges) will be removed if its lifespan is reached.
Some key concepts and parameters are called for at this point:
  • 1.
    Strongly Connected Component (SCC) A maximal set of nodes in which each can reach every other via directed edges. Here, a large SCC indicates a self‐reinforcing community of affordances.
  • 2.
    Lifespan ($L_{S}$). Maximum number of steps a node may survive before it is culled for (i) having no in-edges or (ii) a non-positive net balance of help minus harm.
  • 3.
    Bind-range ($p_{\max}$). Upper bound on the allowable trait-difference for an edge to form; scaled Poisson noise introduces pair-specific variability. This is explained in more detail below.
  • 4.
    Help/Harm ratio ($\rho$). Scales bind-range so that overall fraction of potential interactions are harmful are $\rho$; $(1-\rho)$ are helpful.
  • 5.
    Node traits. Each new node draws four independent traits: $h_{\text{in}},h_{\text{out}}$ or ”help in” and ”help out”, and $\gamma_{\text{in}},\gamma_{\text{out}}$ or ”harm in” and ”harm out.”
    $(h_{\text{in}},h_{\text{out}},\gamma_{\text{in}},\gamma_{\text{out}})\in[0,1]^{4}$ and govern help/harm matching.
  • 6.
    Removal threshold ($\eta$). A node is deleted once $\frac{\text{harm}_{\text{in}}}{(\text{help}_{\text{in}}+\text{harm}_{\text{in}})}>\eta$. We use $\eta=0.5$ unless stated otherwise.
2.1.1 $p_{\max}$ parameter
We set an upper‐bound parameter $p_{\max}$ (starting at $p_{\max}=0.001$ but varies in some simulations) that represents the largest possible difference in pairwise “catalytic potential” between any two nodes. For pair-specific variability, at the time a node $i$ is added, each ordered pair of nodes $(i,j)$ we draw an integer $k_{ij}\sim\mathrm{Poisson}(\lambda)$ and let the effective probability that a directed edge $i\!\to\!j$ and/or $j\!\to\!i$ appears be
$P_{ij}=\min\!\bigl(1,\;k_{ij}\,p_{\max}\bigr).$
To clarify, this $P_{ij}$ is set once between each entering node and every existing node. In ecological terms $k_{ij}$ could be read as the (random) count of events that would enable $i$ to facilitate or provide affordances to $j$ within the current time‐step, while $p_{\max}$ translates that count into a probability of edge-formation.
@@ 2.2 Eigen Analysis of Adjacency Matrix
At each simulation tick, we generate an adjacency matrix by flattening the MultiGraph to a weighted graph by summing the number of edges: +1 for help edges and -1 for harm edges. For example entry $A_{ij}$ will hold +3 if there are 5 helpful and 2 harmful edges from $i\to j$. Recall that at most 6 edges of one type (help or harm) can form between two nodes. We then approximate the dominant eigenpair (eigenvector and corresponding eigenvalue) $(\hat{\lambda}_{1},\hat{v}_{1})$ using a fixed-iteration power method with normalization at each step and a Rayleigh-quotient estimate,
$\hat{\lambda}_{1}\;=\;\frac{\hat{v}_{1}^{\top}A_{t}\hat{v}_{1}}{\hat{v}_{1}^{\top}\hat{v}_{1}},$
@@
$B_{t}\;=\;A_{t}\;-\;\hat{\lambda}_{1}\,\hat{v}_{1}\hat{v}_{1}^{\top},$
@@
The principal eigenvector for the adjacency matrix encodes the principal or dominant direction of flows or activity in the network. It’s corresponding principal eigenvalue gives a measure to the ’strength’ of that dominance. An increasing steady principal eigenvalue also means that average connectivity (and so the potential for feedback/cooperation or harm) is still growing – the graph hasn’t saturated yet.
Our second measure,the spectral gap, is the difference between the two leading eigenvalues: $|\lambda_{1}|-|\lambda_{2}|$. A large and growing spectral gap implies there is one cohesive, resilient strongly connected component. In other words, the principal direction of connectivity has ’pulled away’ from any secondary directions. A small or falling gap would suggest there are many communities and more potential for fragmentation. Put another way, $\lambda_{1}$ tells us how dense the connections are, the spectral gap tells us how internally cohesive that density is.
@@
[…]
@@
    Spectral angle = measures structural change between network states.
@@ 3 Experiments & Results
In a systematic way, we explore how shifting the parameters (lifespan, harm-to-help ratio, binding chance) influence the emergence and persistence of strongly connected components (SCCs) in our directed network. Our primary interest was the evolution and robustness of cooperative structures (large SCCs) in networks characterized by varying degrees of antagonism (harm-to-help ratio). Eigen Analysis was performed to give a higher level characterization of the network’s behavior short and long run behavior.
@@
    Absolute Mean Dominant Eigenalue: Removing the sign of the above reveals just the intensity of any shifts in the principal eigenvector, positive or negative.
@@
    Spectral Gap: The difference between the first and second principal eigenvalues: $|\lambda_{1}|-|\lambda_{2}|$.
@@ 3.1 Harm-To-Help Ratio and Lifespan Dynamics
Figures 4,5,6 show the results of the simulation from solely increasing the Harm-To-Help ratio from $\rho=0.3,0.5,0.7$ with all other parameters fixed. At first, we can restrict our view to the differences in the number of nodes and edges at step $t$, the largest SCC size and largest SCC fraction.
@@ 4 Discussion
4.1 Robustness of Cooperation under High Antagonism
Our model demonstrates robustness of cooperative structures (large SCCs) even when antagonistic interactions (harmful edges) significantly outnumber supportive interactions. This finding could be an extension of the concept of ‘islands of cooperation’ in [21]. Specifically, even with high harm-to-help ratios ($\rho>0.5$), SCCs consistently emerge and persist at steady-state equilibria. This implies the presence of inherent resilience within the network structure, which allows it to dynamically reconfigure to maintain cooperative viability despite continuous antagonistic pressure. The results on node “age” vs degree distribution lend another perspective on the long run dynamics, which we discuss later.
@@ 4.1 Robustness of Cooperation under High Antagonism
4.2 Critical Thresholds and Phase Transitions
A central outcome of our simulations is the emergence of a phase transition in the system behavior at the harm-to-help ratio $\rho\approx 0.6$. (Figure 9, 10) For values of $\rho$ below this threshold, the strongly connected component (SCC) grows steadily with the total node population, reflecting a regime of sustained interaction and widespread mutual support. Above this threshold, however, SCC growth transitions from indefinite expansion to a bounded, fluctuating steady state.
This transition is evident in the late-time population growth rate (measured over the final 500 steps), which decreases sharply and crosses zero as a function of $\rho$ (Figure 9). A corresponding collapse in the mean network degree is shown in Figure 10, indicating a loss of interaction density that accompanies the change in growth dynamics. Together, these results identify a clear qualitative boundary separating the regimes of population growth from bounded population dynamics.
Figure 9: The late-time (last 500 steps) population growth rate $g$ as a function of the harm-to-help ratio $\rho$. For small $\rho$, the system exhibits sustained positive growth. As $\rho$ increases, the growth rate decreases sharply and crosses zero near $\rho\approx 0.6$, marking a transition from sustained growth to a bounded population regime.
Figure 10: Mean total degree (in + out) as a function of the harm-to-help ratio $\rho$. As $\rho$ increases, the average connectivity of the network decreases sharply, with a pronounced collapse occurring near $\rho\approx 0.6$. This reduction in interaction density coincides with the transition from sustained to bounded population growth observer in the above figure, suggesting the loss of sufficient interaction capacity underlies the change in population dynamics.
While the harm-to-help ratio $\rho$ determines whether the system exhibits sustained or bounded growth, the binding chance $p_{max}$ governs the scale at which global coordination emerges. Figure 11 shows the fraction of nodes in the largest strongly connected component as a function of population size for different values of $p_{max}$. Increasing $p_{max}$ sharply reduces the critical population size required for SCC formation, indicating a transition in coordination onset rather than final network structure. This behavior resembles classical phase transitions in percolation theory and random graph models, where a threshold density of connections is required for a ”giant” connected component to emerge. [8] Together, these results demonstrate that $\rho$ and $p_{max}$ control distinct but complementary aspects of the system’s macroscopic behavior.
@@ 4.5 Binding Range as Ecological Analog
The binding range parameter, which can be though of as the probability or “ease” of forming interactions (akin to biochemical or ecological compatibility), accelerates the dynamics of community structuring. It is interesting this parameter does not radically alter long-term outcomes but modulates the speed and slightly reduces the SCC’s ceiling under antagonistic conditions. In ecological contexts, this could suggest ecosystems with higher interaction potentials (dense signaling, physical proximity, etc) might achieve stable cooperation more rapidly and sustain slightly larger cohesive structures despite competitive pressures. Nevertheless, even in cases with a low binding range, it is “just a matter of time” before cooperative communities (autocatalytic sets, ecosystems (biological, economic, cultural), etc.) will be formed.

### v4

@@ Abstract
We present a random threshold-directed network model [24] that integrates node-specific traits with dynamic edge formation and node removal, simulating arbitrary levels of cooperation and competition. In our framework, intrinsic node values determine directed links through various threshold rules. Our model generates a multi-digraph with signed edges (reflecting support/antagonism, labeled “help”/“harm”), which ultimately yields two parallel yet interdependent threshold graphs. Incorporating temporal growth and node turnover in our approach allows exploration of the evolution, adaptation, and potential collapse of communities and reveals regime changes in both connectivity and resilience.
@@
keywords:
adaptive threshold networks , cooperation under antagonism , strongly connected components , regime change , crossover , collective affordance sets , emergent organization , complex systems biology , microbial communities
@@
{graphicalabstract}[Uncaptioned image]
{highlights}
Even in antagonistic environments, the system finds “order for free”: a mutually enabling SCC self-assembles and persists.
When diversity and binding chance cross a threshold, quantity gives way to quality: a giant SCC takes shape.
@@ 2 Model
2.0.1 Model definition (birth–binding–culling dynamics)
We simulate an evolving signed directed multigraph $G_{t}=(V_{t},E_{t})$ using NetworkX’s MultiDiGraph, where vertices represent agents and directed edges represent directed interactions. Time is discrete, $t=1,2,\dots$, and each time-step consists of: (i) adding one new node, (ii) probabilistically forming directed help/harm edges between the newcomer and all incumbents, (iii) aging all nodes by one tick, and (iv) removing any nodes that become eligible for deletion.
Parameters.
The dynamics are governed by the following global parameters:
  • 1.
    $\lambda>0$: Poisson rate controlling heterogeneity in opportunities for pairwise interaction.
  • 2.
    $p_{\max}\in(0,1]$: baseline upper bound used to convert opportunity into an effective binding threshold/probability.
  • 3.
    $\rho\in[0,1]$: global help–harm mixture; $\rho$ is the probability that a successful binding event is harmful (and $1-\rho$ that it is helpful).
  • 4.
    $m_{\max}\in\mathbb{N}$: maximum number of parallel edges allowed per ordered pair and sign (we use $m_{\max}=6$ unless stated otherwise).
  • 5.
    $L_{S}\in\mathbb{N}$: minimum age before a node becomes eligible for removal (a lifespan/juvenile protection period, $L_{S}=100$ by default).
  • 6.
    $\eta\in[0,1]$: removal threshold for the fraction of incoming harm (we use $\eta=0.5$ unless stated otherwise).
Node traits (intrinsic capacities).
When a new node $i$ enters, it is endowed with four independent traits drawn i.i.d. uniformly from $[0,1]$:
  • 1.
    $h_{\text{in}}(i)$: receptivity to beneficial (help) interactions,
  • 2.
    $h_{\text{out}}(i)$: capacity to provide beneficial (help) interactions,
  • 3.
    $\gamma_{\text{in}}(i)$: vulnerability to detrimental (harm) interactions,
  • 4.
    $\gamma_{\text{out}}(i)$: capacity to inflict detrimental (harm) interactions.
We treat help and harm as independent dimensions rather than opposite ends of a single axis. Biologically, an organism can plausibly both support mutualists and inhibit competitors via distinct pathways, so the ability to help need not constrain the ability to harm (and vice versa).
Pair-specific binding opportunity and effective threshold.
When node $i$ enters at time $t$, it “tests” every incumbent node $j\in V_{t-1}$ in both directions $(i\to j)$ and $(j\to i)$. For each ordered pair $(u,v)\in\{(i,j),(j,i)\}$ we draw
$k_{uv}\sim\mathrm{Poisson}(\lambda),\qquad P_{uv}=\min\bigl(1,\;k_{uv}\,p_{\max}\bigr).$
Here $k_{uv}$ represents the (random) number of opportunities for $u$ to interact with $v$ during this time-step, and $P_{uv}$ converts that opportunity into an effective binding scale capped at $1$.
Edge formation (help vs harm, with multiplicity).
For each ordered pair $(u,v)$ involving the newcomer and an incumbent, we perform up to $m_{\max}$ binding attempts. Each attempt proceeds as follows:
  1. 1.
    Draw the interaction type: harm with probability $\rho$, help with probability $1-\rho$.
  2. 2.
    Conditional on the type, form a directed edge if the relevant trait mismatch is within the effective threshold:
    $\text{help:}\quad|h_{\text{out}}(u)-h_{\text{in}}(v)|<P_{uv},\qquad\text{harm:}\quad|\gamma_{\text{out}}(u)-\gamma_{\text{in}}(v)|<P_{uv}.$
Successful attempts add a parallel directed edge of the corresponding sign; unsuccessful attempts add nothing. This allows repeated affordances or repeated adverse interactions between the same ordered pair while preventing unbounded multiplicity.
Aging and culling (node removal rule).
After binding is evaluated for the new node, every node ages by one tick. A node $x$ becomes eligible for removal only once its age is at least $L_{S}$. Eligible nodes are removed (along with all incident edges) if either:
  1. 1.
    $x$ has no incoming edges (no support or influence from others), or
  2. 2.
    the incoming harm fraction exceeds $\eta$:
    $\frac{\mathrm{harm}_{\mathrm{in}}(x)}{\mathrm{help}_{\mathrm{in}}(x)+\mathrm{harm}_{\mathrm{in}}(x)}>\eta.$
Edges do not decay on their own; they disappear only when one endpoint is deleted. Repeating this birth–binding–culling cycle yields the evolving signed network used in subsequent experiments. We use $\eta=0.5$ unless specified otherwise.
@@ 2.0.1 Model definition (birth–binding–culling dynamics)
AABBCCDDEE
AABBDDEE
Figure 1: Sample simulation showing graph generation and node removal. Top row (left to right): Steps 2–4. Bottom row: Step 5 (E added, C marked for removal) and Step 6 (C and its edges removed). For clarity, the lifespan condition is omitted; in the full model, a node is only eligible for removal once its age is at least $L_{S}$.
2.0.2 Sample Simulation
@@
Step 2: Node B added, harm (red) edge added $B\to A$ as $|\gamma_{\text{in},A}-\gamma_{\text{out},B}|<P_{BA}$, and help (blue) edge added $A\to B$ as $|h_{\text{out},A}-h_{\text{in},B}|<P_{AB}$.
@@
Step 5: Node E added, 3 red links added. Node C now receives 3 harm edges (from B and E) and only 2 help edges (from B and D), so $\frac{3}{5}=0.6>\eta=0.5$. C is marked for removal (assuming its age exceeds $L_{S}$).
Step 6: Node C (and all its incident edges) removed. The network continues with nodes A, B, D, E.
@@ 2.1 Eigen Analysis of Adjacency Matrix
At each simulation tick, we generate an adjacency matrix by flattening the MultiGraph to a weighted graph by summing the number of edges: +1 for help edges and -1 for harm edges. Recall that at most 6 edges of one type (help or harm) can form between two nodes.
For example, in the above Figure 1 sample simulation at Step 5 (nodes A, B, C, D, E), the signed adjacency matrix $\mathbf{A}_{t}$ (where +1+1 denotes a help edge and 1-1 a harm edge, summed over parallel edges) is:
[…]
Entry $\mathbf{A}_{ij}$ indicates the net interaction from node $i$ to node $j$. For instance, $\mathbf{A}_{CA}=+1$ reflects C’s net helpful influence on A (2 help edges minus 1 harm edge).
We then approximate the dominant eigenpair (eigenvector and corresponding eigenvalue) $(\hat{\lambda}_{1},\mathbf{v}_{1})$ using a fixed-iteration power method with normalization at each step and a Rayleigh-quotient estimate,
$\hat{\lambda}_{1}\;=\;\frac{\mathbf{v}_{1}^{\top}\mathbf{A}_{t}\mathbf{v}_{1}}{\mathbf{v}_{1}^{\top}\mathbf{v}_{1}},$
@@
$\mathbf{B}_{t}\;=\;\mathbf{A}_{t}\;-\;\hat{\lambda}_{1}\,\mathbf{v}_{1}\mathbf{v}_{1}^{\top},$
@@
The principal eigenvector for the adjacency matrix encodes the principal or dominant direction of flows or activity in the network. Its corresponding principal eigenvalue gives a measure to the ’strength’ of that dominance. An increasing steady principal eigenvalue also means that average connectivity (and so the potential for feedback/cooperation or harm) is still growing – the graph hasn’t saturated yet.
Our second measure, the spectral gap, is the difference between the two leading eigenvalues: $|\hat{\lambda}_{1}|-|\hat{\lambda}_{2}|$. A large and growing spectral gap implies there is one cohesive, resilient strongly connected component. In other words, the principal direction of connectivity has ’pulled away’ from any secondary directions. A small or falling gap would suggest there are many communities and more potential for fragmentation. Put another way, $\lambda_{1}$ tells us how dense the connections are, the spectral gap tells us how internally cohesive that density is.
@@
$\theta_{t}=\arccos\!\left(\frac{\mathbf{v}_{t-1}^{\top}\mathbf{v}_{t}}{\|\mathbf{v}_{t-1}\|\,\|\mathbf{v}_{t}\|}\right).$
@@
    Principal angle = measures structural change between network states.
@@ 3 Experiments & Results
In a systematic way, we explore how shifting the parameters (lifespan, harm-to-help ratio, binding chance) influence the emergence and persistence of strongly connected components (SCCs) in our directed network. Our primary interest was the evolution and robustness of cooperative structures (large SCCs) in networks characterized by varying degrees of antagonism (harm-to-help ratio). Eigen Analysis was performed to give a higher level characterization of the network’s behavior in the short- and long-run.
@@
    Absolute Mean Dominant Eigenvalue: Removing the sign of the above reveals just the intensity of any shifts in the principal eigenvector, positive or negative.
@@
    Spectral Gap: The difference between the first and second principal eigenvalues: $|\hat{\lambda}_{1}|-|\hat{\lambda}_{2}|$.
@@ 3.1 Harm-To-Help Ratio and Lifespan Dynamics
Figures 3,4,5 show the results of the simulation from solely increasing the Harm-To-Help ratio from $\rho=0.3,0.5,0.7$ with all other parameters fixed. At first, we can restrict our view to the differences in the number of nodes and edges at step $t$, the largest SCC size and largest SCC fraction.
@@ 4 Discussion
4.1 Robustness of SCCs under High Antagonism
Our model demonstrates robustness of large SCC structure even when antagonistic interactions (harmful edges) significantly outnumber supportive interactions. This finding could be an extension of the concept of ‘islands of cooperation’ in [21]. Specifically, even with high harm-to-help ratios ($\rho>0.5$), SCCs consistently emerge and persist at steady-state equilibria. This implies the presence of inherent resilience within the network structure, which allows it to dynamically reconfigure to maintain cooperative viability despite continuous antagonistic pressure. The results on node “age” vs degree distribution lend another perspective on the long run dynamics, which we discuss later.
@@ 4.1 Robustness of SCCs under High Antagonism
4.2 Thresholds and Regime Shifts
A central outcome of our simulations is the emergence of a regime change near the harm-to-help ratio $\rho_{c}\approx 0.6$. (Figure 8, 9) For values of $\rho$ below this threshold, the strongly connected component (SCC) grows steadily with the total node population, reflecting a regime of sustained interaction and widespread mutual support. Above this threshold, however, SCC growth transitions from indefinite expansion to a bounded, fluctuating steady state.
This transition is evident in the late-time population growth rate, which decreases as $\rho$ increases and crosses zero near $\rho_{c}\approx 0.6$ (Figure 8). We define the late-time growth rate $g$ as the average per-step change in population size over the final $W=500$ time steps:
$g\;=\;\frac{|V_{T}|-|V_{T-W}|}{W}.$
A corresponding collapse in the mean network degree is shown in Figure 9, indicating a loss of interaction density that accompanies the change in growth dynamics.
Figure 8: The late-time (last 500 steps) population growth rate $g$ as a function of the harm-to-help ratio $\rho$. For small $\rho$, the system exhibits sustained positive growth. As $\rho$ increases, the growth rate decreases and crosses zero near $\rho_{c}\approx 0.6$, marking a shift between distinct dynamical regimes, from sustained growth to a bounded population regime.
Figure 9: Mean total degree (in + out) as a function of the harm-to-help ratio $\rho$. As $\rho$ increases, the average connectivity of the network decreases, with a marked reduction occurring near $\rho_{c}\approx 0.6$. This reduction in interaction density coincides with the crossover from sustained to bounded population growth observed in Figure 8, suggesting the loss of sufficient interaction capacity underlies the shift between dynamical regimes in population behavior.
While the harm-to-help ratio $\rho$ determines whether the system exhibits sustained or bounded growth, the binding chance $p_{max}$ governs the scale at which global coordination emerges. Figure 10 shows the fraction of nodes in the largest strongly connected component as a function of population size for different values of $p_{max}$. Increasing $p_{max}$ reduces the population scale required for SCC formation, indicating a change in coordination regime onset rather than final network structure.
This behavior is reminiscent of the connectivity thresholds in percolation and random graph models, in which a “giant” component becomes likely once the effective connection density is sufficiently high. [8] In our finite, stochastic setting the onset is rounded, but increasing $p_{max}$ consistently shifts coordination to smaller population sizes. Together, these results indicate that $\rho$ governs the growth–boundedness regime, whereas $p_{max}$ governs the population scale at which coordinated SCC structure typically emerges.
@@ 4.5 Binding Range as Ecological Analog
The binding range parameter, which can be thought of as the probability or “ease” of forming interactions (akin to biochemical or ecological compatibility), accelerates the dynamics of community structuring. It is interesting this parameter does not radically alter long-term outcomes but modulates the speed and slightly reduces the SCC’s ceiling under antagonistic conditions. In ecological contexts, this could suggest ecosystems with higher interaction potentials (dense signaling, physical proximity, etc) might achieve stable cooperation more rapidly and sustain slightly larger cohesive structures despite competitive pressures. Nevertheless, even in cases with a low binding range, it is “just a matter of time” before cooperative communities (autocatalytic sets, ecosystems (biological, economic, cultural), etc.) will be formed.